# Federated Deep Reinforcement Learning for Privacy-Preserving Robotic-Assisted Surgery

### v1

@@ II-A1 Task-Specific Accuracy
$A_{task}=\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{t=1}^{T}\mathbb{I}(a_{t}^{i}=a_{t}^{*})}{T}$ (8)
@@ II-A2 Surgical Risk Mitigation Score
$R_{\text{mit}}=1-\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{t=1}^{T}D_{t}^{i}}{T}$ (9)
@@
This formulation ensures a comprehensive risk assessment by balancing force control, tissue integrity, and error minimization. Since lower risk is preferable, the score is structured as 1 minus the average risk per surgery, making $R_{\text{mit}}$ an increasing metric, where higher values indicate safer policy performance. This method integrates numerous risk factors, allowing for an objective and data-driven assessment of RL-based surgical protocols, thereby facilitating the choice of optimal, risk-conscious strategies for both autonomous and semi-autonomous robotic operations.
@@ II-A3 Dynamic Policy Adaptation Rate
$\alpha_{adapt}=\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{t=1}^{T}\mathbb{I}\left(d(s_{t},s_{t-1})>\theta\right)\cdot\mathbb{I}\left(a_{t}\neq a_{t-1}\right)}{\sum_{t=1}^{T}\mathbb{I}\left(d(s_{t},s_{t-1})>\theta\right)}$ (11)
@@ II-B Algorithmic Clarity and Computational Complexity
The FDRL Algorithm Algo.1 ensures privacy-preserving model training in a FL setting for RAS. The algorithm is structured into three primary stages. In the first stage, each hospital independently trains its RL policy using its private dataset while ensuring privacy through DP noise injection before transmitting model updates. The second stage involves secure federated aggregation, where the Federated Aggregator (FA) collects encrypted policy updates from multiple hospitals and processes them using SMPC and HE to maintain strict privacy compliance. Finally, in the third stage, dynamic policy selection takes place, where the surgical robot evaluates federated policies using predefined surgical performance metrics and selects the optimal policy for RAS.
@@
1:Hospitals =$\mathcal{H}=\{H_{1},H_{2},\dots,H_{N}\}$, datasets $\mathcal{D}_{i}$, local policies $\pi_{i}$, global policy $\pi_{G}$
@@
To address this, we incorporated weighted local updates (Eq. 3), with a regularisation term $\lambda$ that penalises divergence between local models and the global model. Furthermore, we simulated varying levels of non-IID environments by adjusting the heterogeneity factor from 0 (Independent and Identically Distributed (IID)) to 1 (highly non-IID), demonstrating that the proposed FDRL framework consistently maintains a stable accuracy of $>92\%$, even under severe heterogeneity conditions.
@@
$\nabla L_{i}^{\prime}=\nabla L_{i}+\mathcal{N}(0,\sigma^{2})$ (12)
@@
$\epsilon=\frac{\alpha}{2\sigma^{2}}$ (13)
@@
Despite the higher computational overhead introduced by HE compared to standard averaging techniques, the privacy-security tradeoff ensures that patient data confidentiality is preserved without direct exposure. The optimization of policy selection Minimises computational costs, enabling real-time decision-making in RAS. The proposed FDRL framework remains computationally feasible for real-world deployment, striking a balance between privacy preservation and model efficiency. Future research will explore hardware acceleration techniques, such as quantized FL and edge computing integration, to further reduce computational overhead and improve scalability.
@@
We compare our FDRL framework with a centralized RL framework as a baseline (where all policies are pooled together) for privacy effectiveness using the following privacy metrics. Additionally, we introduce an experimental evaluation that systematically analyzes how different privacy settings ($\epsilon$, $\sigma^{2}$) impact surgical performance in terms of accuracy, safety, and adaptability. A privacy utility tradeoff plot is included to evaluate how privacy constraints influence surgical precision, task success rates, and training efficiency.
@@ III-A Privacy Leakage Rate (PLR) Calculation
$\text{PLR}_{FL}=\frac{1}{N}\sum_{i=1}^{N}\frac{I(W_{i};D_{i})}{H_{2}(D_{i})}$ (16)
where $W_{i}$ represents the policy weights at hospital $i$, $D_{i}$ is the local dataset at hospital $i$, and $H_{2}(D_{i})$ denotes Shannon entropy (in bits) for dataset $D_{i}$. For Centralized Training, where data is aggregated across all hospitals, the PLR is given by
$\text{PLR}_{Central}=\frac{I(W_{global};D_{all})}{H_{2}(D_{all})}$ (17)
@@ III-B Policy Divergence Across Hospitals
Policy divergence measures how different local policies are from a globally trained policy, serving as a proxy for privacy
$D_{KL}(\pi_{i}||\pi_{global})=\sum_{s}\sum_{a}\pi_{i}(a|s)\log\frac{\pi_{i}(a|s)}{\pi_{global}(a|s)}$ (18)
@@
$D_{FL}=\frac{1}{N}\sum_{i=1}^{N}D_{KL}(\pi_{i}||\pi_{FL})$ (19)
For centralized training
$D_{Central}=D_{KL}(\pi_{global}||\pi_{centralized})$ (20)
@@ III-C Differential Privacy and Gradient Noise
$g^{\prime}=g+N(0,\sigma^{2})$ (21)
where $g$ is the original gradient and $N(0,\sigma^{2})$ is Gaussian noise. The privacy budget $\epsilon$ is computed using the RDP framework
$\epsilon_{FL}=\frac{\alpha}{2\sigma^{2}}$ (22)
For centralized RL
$\epsilon_{Central}=\frac{\alpha}{2\sigma_{central}^{2}}$ (23)
here, lower $\epsilon$ means better privacy.
@@ III-C1 Overall Privacy Effectiveness (OPE)
$OPE=\lambda_{1}(1-PLR)+\lambda_{2}D_{KL}+\lambda_{3}e^{-\epsilon}$ (24)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are weights based on importance. If $OPE_{FL}>OPE_{Central}$, then FL provides stronger privacy. If $OPE_{FL}<OPE_{Central}$, then centralized training is more private.
@@ III-D Homomorphic Encryption and Secure Aggregation
In the proposed FDRL framework, HE is explicitly employed to safeguard privacy during the aggregation of local model parameters. Each participating hospital encrypts its local model parameters (weights or gradients) using HE before sending them to the federated aggregation server. This process ensures strict privacy of all model updates throughout the entire aggregation procedure. Specifically, HE enables the federated aggregator to perform arithmetic operations (e.g., addition and averaging) directly on encrypted data, thereby preserving privacy by preventing the exposure of sensitive intermediate gradient or weight information. The secure aggregation process using HE operates as follows. Each hospital encrypts its local model parameters using a public key, yielding encrypted parameters $\text{Enc}(\theta_{i})=HE.Enc_{pk}(\theta_{i})$, where $HE.Enc$ denotes the HE function. These encrypted parameters are then securely transmitted to the federated aggregation server, i.e., Hospital $\text{Hospital }I\rightarrow\text{Server}:\text{Enc}(\theta_{i})$.
The federated aggregator performs a weighted aggregation directly on the encrypted data using the additive homomorphic property, computing the global encrypted parameters as $\text{Enc}(\theta_{\text{global}})=\sum_{i=1}^{N}\frac{n_{i}}{n}\text{Enc}(\theta_{i})$. Here, $N$ represents the total number of hospitals, $n_{i}$ is the number of training samples at hospital $i$, and $n$ is the total number of training samples across all hospitals. After aggregation, decryption occurs only at the trusted global server using the private key $sk$, such that $\theta_{\text{global}}=HE.Dec_{sk}\left(\text{Enc}(\theta_{\text{global}})\right)$, where $HE.Dec$ denotes the homomorphic decryption function. Finally, the decrypted global model parameters $\theta_{\text{global}}$ are securely distributed back to all local hospitals, completing the federated training round, i.e., $\text{Server}\rightarrow\text{Hospitals}:\theta_{\text{global}}$.
@@ IV-A Simulation Environment Development
We developed a synthetic surgical environment using Python, leveraging libraries such as NumPy for numerical computations and Matplotlib/Seaborn for data visualization. This environment was designed to simulate surgical procedures across multiple hospital sites, each possessing unique patient data distributions and surgical specializations. The environment models surgical tasks as MDP, which provide a structured framework for representing sequential decision-making problems. In this context, the state space represents patient-specific parameters, including vital signs, medical imaging data, and physiological states. The action space encompasses robotic tool movements, incision strategies, and other surgical interventions. The reward function is designed to incentivize optimal surgical outcomes, penalizing errors and rewarding precision. Transition probabilities model the dynamic changes in the patient’s state based on the actions taken by the surgical robot. This comprehensive simulation environment allowed us to thoroughly evaluate the FDRL framework under various realistic conditions.
Figure 3: Performance Comparison of Federated versus Centralised RL. (a) PLR over 50 global rounds, where lower PLR for FDRL indicates better privacy preservation compared to the Centralised approach. (b) Average Kullback–Leibler (KL) divergence between local and aggregated (global) policy distributions in FDRL, demonstrating the extent of divergence among local models during training. (c) OPE over global rounds, computed as a weighted combination of (1 – PLR), KL divergence, and an effective DP decay term, highlighting the superior privacy–performance trade-off achieved by Federated RL relative to Centralised RL
@@ IV-B Differential Privacy Parameters and Ablation Study
We considered $\epsilon=1$ as the default privacy budget in DP, aligning with healthcare privacy standards, while varying $\sigma^{2}$ from $0.01$ to $1$. The choice of $\epsilon$ balances privacy preservation with acceptable model performance, as recommended in prior healthcare FL studies. Furthermore, we conducted an ablation study to isolate the privacy-preserving components in our proposed FDRL framework. The PLR and accuracy trade-offs across these settings, confirming that integrating DP and HE significantly reduces privacy leakage by approximately $60\%$, albeit with marginal computational overhead and negligible accuracy loss of approximately $1.5\%$. Synthetic datasets were generated to emulate diverse surgical scenarios, ensuring a broad range of patient conditions and surgical complexities. Each hospital’s dataset was created with varying degrees of heterogeneity, simulating real-world differences in patient demographics and surgical practices. The data included simulated medical images, vital signs, and surgical history, allowing for a comprehensive evaluation of the FDRL framework’s performance and privacy characteristics. The generation of synthetic data enabled us to control and manipulate variables such as data distribution and heterogeneity, providing a robust testbed for our experiments. While HE and Secure Aggregation ensure strong privacy guarantees, their computational overhead is non-trivial, particularly in real-time RAS. Specifically, encryption and decryption operations add an average latency of 0.7 seconds per communication round in our simulations. To mitigate this, lightweight encryption schemes (e.g., partially homomorphic schemes or hybrid models combining symmetric cryptography for non-sensitive data) are proposed for future deployment.
@@
The simulations were executed on a high-performance computing cluster equipped with multi-core Intel Xeon processors and NVIDIA GPUs for accelerated deep-learning computations. High-speed network connectivity was utilized to simulate federated communication between hospital sites. Each node was equipped with 32GB of RAM, ensuring efficient execution of the simulations and accurate evaluation of the framework’s performance. This robust hardware setup allowed us to conduct extensive experiments and analyze the results with high precision. The simulations were designed to reflect realistic surgical scenarios and evaluate the framework’s performance under diverse conditions. Three hospitals participated, each with a dataset of 100 samples. The state and action dimensions were set to five and three, respectively. FL ran for 50 rounds, each with five local epochs, while Centralized Learning (CL) lasted 15 epochs.DP noise standard deviation varied from 0.01 to 1 for FL and was fixed at 0.1 for CL. The heterogeneity factor, controlling dataset variation across hospitals, ranged from 0 to 1. The OPE metric’s weighting coefficients were set to $\lambda_{1}=0.3$, $\lambda_{2}=0.4$, and $\lambda_{3}=0.3$, ensuring a comprehensive analysis of the framework’s behaviour. For Fig. 3, 2(a) compares PLR in Federated and Centralized RL, showing that FL significantly reduces PLR, indicating stronger privacy preservation by decentralizing model training and avoiding direct data sharing. The higher PLR in CL highlights the risk of information leakage due to data aggregation. Fig. 2(b) presents the Kullback-Leibler Divergence (KL) divergence between locally trained and global policies. The higher divergence in Federated RL suggests greater policy variation across hospitals, enhancing privacy by reducing the risk of dataset reconstruction from the global model. Fig. 2(c) shows the OPE score, confirming Federated RL ’s superior privacy-preserving capabilities by integrating PLR, policy divergence, and DP constraints. The results across 50 global rounds demonstrate the stability and effectiveness of FL in balancing privacy and utility.
@@
Figure 5: Comparison of FL and CL across key evaluation metrics: Task-Specific Accuracy, Surgical Risk Mitigation, PLR, and Overall Privacy Effectiveness (OPE).
@@
Fig. 4 presents the impact of DP noise standard deviation on both Accuracy and Privacy Effectiveness in a FL environment. The x-axis represents increasing levels of noise added for DP, ranging from 0.01 to 1.00, effectively capturing the spectrum of privacy protection strength. The heatmap highlights the inverse relationship between these two metrics as the noise standard deviation increases, Privacy effectiveness improves, indicated by a shift towards warmer colours, while accuracy declines, reflected by a transition towards cooler colours. This visualization effectively demonstrates the privacy-utility trade-off inherent in differentially private FL higher noise ensures stronger privacy guarantees but comes at the cost of reduced model accuracy, and vice versa. Fig. 5 demonstrates the performance differences between FL and CL across four key evaluation metrics in privacy-preserving RAS. The selected metrics Task-Specific Accuracy, Surgical Risk Mitigation, PLR, and OPE offer a comprehensive assessment of both approaches.
@@
FL outperforms CL in key privacy and security aspects while maintaining comparable task accuracy. Surgical Risk Mitigation, which measures the ability to Minimise errors and improve procedural safety, is 15.2% higher in FL than in CL. This indicates that FL-trained models adapt more effectively to dynamic surgical environments, potentially reducing risks during real-world deployment. Additionally, PLR is reduced by 60% in FL, highlighting its advantage in securing sensitive patient data. Unlike CL, which requires direct data aggregation and exposes information to central repositories, FL performs decentralized learning, inherently enhancing privacy preservation.
@@
Fig. 6 shows the interplay between data heterogeneity, privacy-preserving noise, and model accuracy in a FL setting. The 3D surface trend reveals a clear inverse correlation: model accuracy declines as either data heterogeneity or noise standard deviation increases. This aligns with existing FL literature, where data divergence across clients impairs global model convergence and generalization. Simultaneously, higher noise levels while improving privacy further reduce accuracy, reflecting the classic privacy-utility trade-off.
@@
Figure 7: Comparative Analysis of Federated Learning and Centralized Learning Across Various Performance Metrics.
Fig. 7 provides a comparison of FL and CL across various performance metrics. The results demonstrate that FL, despite its privacy-preserving advantages, achieves comparable or even superior accuracy in certain aspects. Specifically, FL exhibits a 15% increase in risk mitigation compared to CL, underscoring its potential for enhancing safety and reliability. Furthermore, FL significantly outperform CL in terms of privacy, with P-Leakage values reduced by 60% and Policy Divergence reduced by 50%. The OPE score, combining both P-Leakage and policy divergence, is also higher for FL, indicating a more favourable balance between privacy and performance. These results underscore the potential of FL to match or surpass the accuracy of CL while offering stronger privacy assurances, especially in sensitive areas like healthcare.
In traditional machine learning, the IID assumption is key for model convergence and generalization. Nonetheless, this assumption often fails in real-world medical FL, particularly in RAS, due to substantial variability in hospital data stemming from differences in patient demographics, disease prevalence, genetic profiles, clinical protocols, and data annotation methods. These inherent non-IID traits typically lead to delayed global model convergence, client drift, reduced generalizability, and increased communication overhead. To address these difficulties, our proposed FDRL framework incorporates: (i) weighted aggregation to handle data imbalance, (ii) proximal regularization, drawing on FedProx, to reduce client drift, (iii) a MSS for dynamic policy adaptation tailored to personalized surgical decisions, and (iv) simulated non-IID settings for robustness testing. Experimental findings reveal that even with a high degree of data heterogeneity ($H=0.8$), our FDRL framework maintains stable surgical accuracy ( 91%), Minimises policy divergence, and achieves an optimal privacy-utility balance, confirming its efficacy in highly varied medical environments.

### v2

@@ II-A1 Task-Specific Accuracy
$A_{task}=\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{t=1}^{T}\mathbb{I}(a_{t}^{i}=a_{t}^{*})}{T}.$ (8)
@@ II-A2 Surgical Risk Mitigation Score
$R_{\text{mit}}=1-\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{t=1}^{T}D_{t}^{i}}{T}.$ (9)
@@
This formulation ensures a comprehensive risk assessment by balancing force control, tissue integrity, and error minimisation. Since lower risk is preferable, the score is structured as 1 minus the average risk per surgery, making $R_{\text{mit}}$ an increasing metric, where higher values indicate safer policy performance. This method integrates numerous risk factors, allowing for an objective and data-driven assessment of RL-based surgical protocols, thereby facilitating the choice of optimal, risk-conscious strategies for both autonomous and semi-autonomous robotic operations.
@@ II-A3 Dynamic Policy Adaptation Rate
$\alpha_{adapt}=\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{t=1}^{T}\mathbb{I}\left(d(s_{t},s_{t-1})>\theta\right)\cdot\mathbb{I}\left(a_{t}\neq a_{t-1}\right)}{\sum_{t=1}^{T}\mathbb{I}\left(d(s_{t},s_{t-1})>\theta\right)}.$ (11)
@@ II-B Algorithmic Clarity and Computational Complexity
The FDRL Algorithm 1 ensures privacy-preserving model training in a FL setting for RAS. The algorithm is structured into three primary stages. In the first stage, each hospital independently trains its RL policy using its private dataset while ensuring privacy through DP noise injection before transmitting model updates. The second stage involves secure federated aggregation, where the Federated Aggregator (FA) collects encrypted policy updates from multiple hospitals and processes them using SMPC and HE to maintain strict privacy compliance. Finally, in the third stage, dynamic policy selection takes place, where the surgical robot evaluates federated policies using predefined surgical performance metrics and selects the optimal policy for RAS.
@@
1:Hospitals $\mathcal{H}=\{H_{1},H_{2},\dots,H_{N}\}$, datasets $\mathcal{D}_{i}$, local policies $\pi_{i}$, global policy $\pi_{G}$
@@
To address this, we incorporated weighted local updates (Eq. 3), with a regularisation term $\lambda$ that penalises divergence between local models and the global model. Furthermore, we simulated varying levels of non-IID environments by adjusting the heterogeneity factor from 0 (Independent and Identically Distributed (IID)) to 1 (highly non-IID), demonstrating that the proposed FDRL framework consistently maintains a stable accuracy of $92\%$, even under severe heterogeneity conditions.
@@
$\nabla L_{i}^{\prime}=\nabla L_{i}+\mathcal{N}(0,\sigma^{2}).$ (12)
@@
$\epsilon=\frac{\alpha}{2\sigma^{2}}.$ (13)
@@
Despite the higher computational overhead introduced by HE compared to standard averaging techniques, the privacy-security tradeoff ensures that patient data confidentiality is preserved without direct exposure. The optimisation of policy selection minimises computational costs, enabling real-time decision-making in RAS. The proposed FDRL framework remains computationally feasible for real-world deployment, striking a balance between privacy-preservation and model efficiency. Future research will explore hardware acceleration techniques, such as quantised FL and edge computing integration, to further reduce computational overhead and improve scalability.
@@
We compare our FDRL framework with a centralised RL framework as a baseline (where all policies are pooled together) for privacy effectiveness using the following privacy metrics. Additionally, we introduce an experimental evaluation that systematically analyses how different privacy settings ($\epsilon$, $\sigma^{2}$) impact surgical performance in terms of accuracy, safety, and adaptability. A privacy utility tradeoff plot is included to evaluate how privacy constraints influence surgical precision, task success rates, and training efficiency.
@@ III-A Privacy Leakage Rate (PLR) Calculation
$\text{PLR}_{FL}=\frac{1}{N}\sum_{i=1}^{N}\frac{I(W_{i};D_{i})}{H_{2}(D_{i})}.$ (16)
where $W_{i}$ represents the policy weights at hospital $i$, $D_{i}$ is the local dataset at hospital $i$, and $H_{2}(D_{i})$ denotes Shannon entropy (in bits) for dataset $D_{i}$. For Centralised Training, where data is aggregated across all hospitals, the PLR is given by
$\text{PLR}_{Central}=\frac{I(W_{global};D_{all})}{H_{2}(D_{all})}.$ (17)
@@ III-B Policy Divergence Across Hospitals
Policy divergence measures how different local policies are from a globally trained policy, serving as a proxy for privacy.
$D_{KL}(\pi_{i}||\pi_{global})=\sum_{s}\sum_{a}\pi_{i}(a|s)\log\frac{\pi_{i}(a|s)}{\pi_{global}(a|s)}.$ (18)
@@
$D_{FL}=\frac{1}{N}\sum_{i=1}^{N}D_{KL}(\pi_{i}||\pi_{FL}).$ (19)
For centralised training
$D_{Central}=D_{KL}(\pi_{global}||\pi_{centralized}).$ (20)
@@ III-C Differential Privacy and Gradient Noise
$g^{\prime}=g+N(0,\sigma^{2}).$ (21)
where $g$ is the original gradient and $N(0,\sigma^{2})$ is Gaussian noise. The privacy budget $\epsilon$ is computed using the Rényi Differential Privacy (RDP) framework
$\epsilon_{FL}=\frac{\alpha}{2\sigma^{2}}.$ (22)
For centralised RL
$\epsilon_{Central}=\frac{\alpha}{2\sigma_{central}^{2}}.$ (23)
Lower $\epsilon$ means better privacy.
@@ III-C1 Overall Privacy Effectiveness (OPE)
$OPE=\lambda_{1}(1-PLR)+\lambda_{2}D_{KL}+\lambda_{3}e^{-\epsilon}.$ (24)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are weights based on importance. If $OPE_{FL}>OPE_{Central}$, then FL provides stronger privacy. If $OPE_{FL}<OPE_{Central}$, then centralised training is more private.
@@ III-D Homomorphic Encryption and Secure Aggregation
In the proposed FDRL framework, HE is explicitly employed to safeguard privacy during the aggregation of local model parameters. Each participating hospital encrypts its local model parameters (weights or gradients) using HE before sending them to the federated aggregation server. This process ensures strict privacy of all model updates throughout the entire aggregation procedure. Specifically, HE enables the federated aggregator to perform arithmetic operations (e.g., addition and averaging) directly on encrypted data, thereby preserving privacy by preventing the exposure of sensitive intermediate gradient or weight information. The secure aggregation process using HE operates as follows. Each hospital encrypts its local model parameters using a public key, yielding encrypted parameters $\text{Enc}(\theta_{i})=HE.Enc_{pk}(\theta_{i})$, where $HE.Enc$ denotes the HE function. These encrypted parameters are then securely transmitted to the federated aggregation server, i.e., $\text{Hospital }I\rightarrow\text{Server}:\text{Enc}(\theta_{i})$.
The federated aggregator performs a weighted aggregation directly on the encrypted data using the additive homomorphic property, computing the global encrypted parameters as $\text{Enc}(\theta_{\text{global}})=\sum_{i=1}^{N}\frac{n_{i}}{n}\text{Enc}(\theta_{i})$. Here, $N$ represents the total number of hospitals, $n_{i}$ is the number of training samples at hospital $i$, and $n$ is the total number of training samples across all hospitals. After aggregation, decryption occurs only at the trusted global server using the private key $sk$, such that $\theta_{\text{global}}=HE.Dec_{sk}\left(\text{Enc}(\theta_{\text{global}})\right)$, where $HE.Dec$ denotes the homomorphic decryption function. Finally, the decrypted global model parameters $\theta_{\text{global}}$ are securely distributed back to all local hospitals, completing the federated training round, i.e., $\text{Server}\rightarrow\text{Hospitals}:\theta_{\text{global}}$. The proposed framework assumes a semi-honest setting where aggregation servers and clients may infer sensitive information without deviating from the protocol. The main threats include: (1) model inversion, reconstructing data from gradients; (2) membership inference, detecting training membership; (3) gradient leakage, exposing input-label pairs; and (4) poisoning, biasing global updates. Our use of DP and HE counters these via noise-injected gradients and encrypted aggregation.
@@ IV-A Simulation Environment Development
We developed a synthetic surgical environment using Python, leveraging libraries such as NumPy for numerical computations and Matplotlib/Seaborn for data visualisation. This environment was designed to simulate surgical procedures across multiple hospital sites, each possessing unique patient data distributions and surgical specialisations. The environment models surgical tasks as MDP, which provides a structured framework for representing sequential decision-making problems. In this context, the state space represents patient-specific parameters, including vital signs, medical imaging data, and physiological states. The action space encompasses robotic tool movements, incision strategies, and other surgical interventions. The reward function is designed to incentivise optimal surgical outcomes, penalising errors and rewarding precision. Transition probabilities model the dynamic changes in the patient’s state based on the actions taken by the surgical robot. This comprehensive simulation environment allowed us to thoroughly evaluate the FDRL framework under various realistic conditions.
Figure 3: Performance comparison of Federated versus Centralised RL: (a) PLR over 50 global rounds; (b) KL divergence between local and global policy distributions; (c) OPE as a weighted combination of PLR, KL divergence, and DP decay, demonstrating superior privacy–utility trade-off in Federated RL.
@@ IV-B Differential Privacy Parameters and Ablation Study
We considered $\epsilon=1$ as the default privacy budget in DP, aligning with healthcare privacy standards, while varying $\sigma^{2}$ from $0.01$ to $1$. The choice of $\epsilon$ balances privacy-preservation with acceptable model performance, as recommended in prior healthcare FL studies. Furthermore, we conducted an ablation study to isolate the privacy-preserving components in our proposed FDRL framework. The PLR and accuracy trade-offs across these settings confirm that integrating DP and HE significantly reduces privacy leakage by approximately $60\%$, albeit with marginal computational overhead and negligible accuracy loss of approximately $1.5\%$. Synthetic datasets were generated to emulate diverse surgical scenarios, ensuring a broad range of patient conditions and surgical complexities. Each hospital’s dataset was created with varying degrees of heterogeneity, simulating real-world differences in patient demographics and surgical practices. The data included simulated medical images, vital signs, and surgical history, allowing for a comprehensive evaluation of the FDRL framework’s performance and privacy characteristics. The generation of synthetic data enabled us to control and manipulate variables such as data distribution and heterogeneity, providing a robust testbed for our experiments. While HE and Secure Aggregation ensure strong privacy guarantees, their computational overhead is non-trivial, particularly in real-time RAS. Specifically, encryption and decryption operations add an average latency of 0.7 seconds per communication round in our simulations. To mitigate this, lightweight encryption schemes (e.g., partially homomorphic schemes or hybrid models combining symmetric cryptography for non-sensitive data) are proposed for future deployment.
@@
The simulations were executed on a high-performance computing cluster equipped with multi-core Intel Xeon processors and NVIDIA GPUs for accelerated deep-learning computations. High-speed network connectivity was utilized to simulate federated communication between hospital sites. Each node was equipped with 32GB of RAM, ensuring efficient execution of the simulations and accurate evaluation of the framework’s performance. This robust hardware setup allowed us to conduct extensive experiments and analyse the results with high precision. The simulations were designed to reflect realistic surgical scenarios and evaluate the framework’s performance under diverse conditions. Three hospitals participated, each with a dataset of 100 samples. The state and action dimensions were set to five and three, respectively. FL ran for 50 rounds, each with five local epochs, while Centralised Learning (CL) lasted 15 epochs. DP noise standard deviation varied from 0.01 to 1 for FL and was fixed at 0.1 for CL. The heterogeneity factor, controlling dataset variation across hospitals, ranged from 0 to 1. The OPE metric’s weighting coefficients were set to $\lambda_{1}=0.3$, $\lambda_{2}=0.4$, and $\lambda_{3}=0.3$, ensuring a comprehensive analysis of the framework’s behaviour. For Fig. 3a compares PLR in Federated and centralised RL, showing that FL significantly reduces PLR, indicating stronger privacy-preservation by decentralising model training and avoiding direct data sharing. The higher PLR in CL highlights the risk of information leakage due to data aggregation. Fig. 3b presents the Kullback-Leibler Divergence (KL) between locally trained and global policies. The higher divergence in Federated RL suggests greater policy variation across hospitals, enhancing privacy by reducing the risk of dataset reconstruction from the global model. Fig. 3c shows the OPE score, confirming Federated RL ’s superior privacy-preserving capabilities by integrating PLR, policy divergence, and DP constraints. The results across 50 global rounds demonstrate the stability and effectiveness of FL in balancing privacy and utility.
@@
Figure 5: Comparison of FL and CL across key evaluation metrics: Task-Specific Accuracy, Surgical Risk Mitigation, PLR and Overall Privacy Effectiveness (OPE).
@@
Fig. 4 presents the impact of DP noise standard deviation on both accuracy and privacy effectiveness in a FL environment. The x-axis represents increasing levels of noise added for DP, ranging from 0.01 to 1.00, effectively capturing the spectrum of privacy protection strength. The heatmap highlights the inverse relationship between these two metrics as the noise standard deviation increases. Privacy effectiveness improves, indicated by a shift towards warmer colours, while accuracy declines, reflected by a transition towards cooler colours. This visualization effectively demonstrates the privacy-utility trade-off inherent in differentially private FL higher noise ensures stronger privacy guarantees but comes at the cost of reduced model accuracy, and vice versa. Fig. 5 demonstrates the performance differences between FL and CL across four key evaluation metrics in privacy-preserving RAS. The selected metrics Task-Specific Accuracy, Surgical Risk Mitigation, PLR, and OPE offer a comprehensive assessment of both approaches.
@@
FL outperforms CL in key privacy and security aspects while maintaining comparable task accuracy. Surgical Risk Mitigation, which measures the ability to minimise errors and improve procedural safety, is 15.2% higher in FL than in CL. This indicates that FL-trained models adapt more effectively to dynamic surgical environments, potentially reducing risks during real-world deployment. Additionally, PLR is reduced by 60% in FL, highlighting its advantage in securing sensitive patient data. Unlike CL, which requires direct data aggregation and exposes information to central repositories, FL performs decentralised learning, inherently enhancing privacy-preservation.
@@
Fig. 6 shows the interplay between data heterogeneity, privacy-preserving noise, and model accuracy in a FL setting. The 3D surface trend reveals a clear inverse correlation. model accuracy declines as either data heterogeneity or noise standard deviation increases. This aligns with existing FL literature, where data divergence across clients impairs global model convergence and generalisation. Simultaneously, higher noise levels while improving privacy further reduce accuracy, reflecting the classic privacy-utility trade-off.
@@
Figure 7: Comparative Analysis of Federated Learning and Centralised Learning Across Various Performance Metrics.
Fig. 7 provides a comparison of FL and CL across various performance metrics. The results demonstrate that FL, despite its privacy-preserving advantages, achieves comparable or even superior accuracy in certain aspects. Specifically, FL exhibits a 15% increase in risk mitigation compared to CL, underscoring its potential for enhancing safety and reliability. Furthermore, FL significantly outperforms CL in terms of privacy, with P-Leakage values reduced by 60% and Policy Divergence reduced by 50%. The OPE score, combining both P-Leakage and policy divergence, is also higher for FL, indicating a more favourable balance between privacy and performance. These results underscore the potential of FL to match or surpass the accuracy of CL while offering stronger privacy assurances, especially in sensitive areas like healthcare.
In traditional machine learning, the IID assumption is key for model convergence and generalisation. Nonetheless, this assumption often fails in real-world medical FL, particularly in RAS, due to substantial variability in hospital data stemming from differences in patient demographics, disease prevalence, genetic profiles, clinical protocols, and data annotation methods. These inherent non-IID traits typically lead to delayed global model convergence, client drift, reduced generalisability, and increased communication overhead. To address these difficulties, our proposed FDRL framework incorporates: (i) weighted aggregation to handle data imbalance, (ii) proximal regularization, drawing on FedProx, to reduce client drift, (iii) a MSS for dynamic policy adaptation tailored to personalised surgical decisions, and (iv) simulated non-IID settings for robustness testing. Experimental findings reveal that even with a high degree of data heterogeneity ($H=0.8$), our FDRL framework maintains stable surgical accuracy ( 91%), minimizes policy divergence, and achieves an optimal privacy-utility balance, confirming its efficacy in highly varied medical environments.